### v1

@@ Abstract
Gen-Swarms is an innovative method that leverages and combines the capabilities of deep generative models with reactive navigation algorithms to automate the creation of drone shows. dvancements in deep generative models, particularly diffusion models, have demonstrated remarkable effectiveness in generating high-quality 2D images. Building on this success, various works have extended diffusion models to 3D point cloud generation. In contrast, alternative generative models such as flow matching have been proposed, offering a simple and intuitive transition from noise to meaningful outputs. However, the application of flow matching models to 3D point cloud generation remains largely unexplored. Gen-Swarms adapts these models to automatically generate drone shows. Existing 3D point cloud generative models create point trajectories which are impractical for drone swarms. In contrast, our method not only generates accurate 3D shapes but also guides the swarm motion, producing smooth trajectories and accounting for potential collisions through a reactive navigation algorithm incorporated into the sampling process. For example, when given a text category like Airplane, Gen-Swarms can rapidly and continuously generate numerous variations of 3D airplane shapes. Our experiments demonstrate that this approach is particularly well-suited for drone shows, providing feasible trajectories, creating representative final shapes, and significantly enhancing the overall performance of drone show generation.
@@ 3.2 DDPMs and CNFs
[…]
@@ 3.3 Flow matching for generative modeling
$p_{t}(\mathbf{x}|\mathbf{x}_{0}):=\mathcal{N}\left(\mathbf{x}|t^{\prime}\mathbf{x}_{0},\sigma_{t}\mathbf{I}\right),\hskip 5.69054pt\sigma_{t}=1-(1-\sigma_{\min})t^{\prime},\hskip 5.69054ptt^{\prime}=(T-t)/T,$ (4)
@@ 4.1 Method
$p_{t}(\mathbf{X}|\mathbf{X}_{0}):=\mathcal{N}\left(\mathbf{X}|t^{\prime}\mathbf{X}_{0},\sigma_{t}\mathbf{I}\right),\hskip 5.69054pt\sigma_{t}=1-(1-\sigma_{\min})t^{\prime},\hskip 5.69054ptt^{\prime}=(T-t)/T.$ (8)
@@
$q(\mathbf{z})=q_{\mathbf{w}}\left(\mathbf{w}\right)\bigg{|}\frac{\partial F_{\alpha}}{\partial\mathbf{w}}\bigg{|}^{-1},\hskip 8.53581pt\mathbf{w}=F^{-1}_{\alpha}(\mathbf{z}).$ (10)
@@
$\mathbf{X}_{t-\Delta t}=\mathbf{X}_{t}+\rho(\mathbf{V}^{t}_{\theta},\mathbf{X}_{t},\kappa),\hskip 11.38109pt\forall t\in\{T,T-\Delta t,\ldots,\Delta t\}.$ (13)
@@ 5.2 Experimental Setup
Regarding architecture details, our Gated Contextual Network consists of $6$ layers, while the bijector $F_{\alpha}$ comprises 14 coupling layers. These neural networks, along with the PointNet encoder with embedding size $d=256$, are trained using the Adam optimizer with a linear scheduler for the learning rate. The experiments were conducted on an Intel Core™ i7-12700K processor with 20 cores and an NVIDIA GeForce RTX 3090 GPU. Furthermore, focusing on method hyperparameters, we set $T=1$ and vary the $\Delta t$ rate according to the experiment to ensure a fair comparison. Additionally, we set a minimum security distance between points of $\kappa=2$ meters in our final scale, which results in $\kappa=0.06$ meters at the training scale. This distance is based on the assumption that drones are 0.5 meters in size (0.25m of radius) with a security distance of 1.5 meters between drones. Lastly, we choose ORCA [30] as our reactive navigation method $\rho$.
@@
  • CMF+ORCA: This third baseline is designed to assess the performance of Gen-Swarms sampling. This approach takes the final point cloud generated by CMF and uses directly the ORCA reactive navigation method to generate all trajectories from the random initial point cloud towards the goal.
@@ 5.3 Quantitative results
For each method, we generate 607 samples, matching the test set size to obtain accurate quality metrics for 3D shapes. Additionally, to ensure a fair comparison, we perform sampling over 100 steps ($\Delta t=0.01s.$) for all methods. We compare the performance of Gen-Swarms against the baselines in the Airplane category using the metrics outlined in Sec. 5.2. The results, presented in a real scale of $200\times 200\times 200$ meters, are summarized in Table 1. As observed, Diffusion leads in quality but falls short in collision avoidance, smoothness and energy consumption. Conditional Matching Flow improves the smoothness and provides more direct, lower-energy trajectories, but still suffers from collisions. When leveraging ORCA to generate safe trajectories towards the final 3D point clouds generated by CMF, it fails to reach these positions within the same number of steps, and does not form the final desired shape, as graphically visualized in the fourth column of Fig. 5. In contrast, Gen-Swarms delivers balanced performance, excelling in collision avoidance (collisions are only present in the initial random point cloud, which is why the TRAJ metric is not zero), energy efficiency, and maintaining high quality while ensuring smooth trajectories. This is because Gen-Swarms uses CFM as a closed-loop controller with ORCA as an additional reactive navigation layer inside the loop.
@@ 5.4 Qualitative results
We illustrate in Fig. 5 an example of a final 3D shape produced by each method in the Airplane category. Drones highlighted in red would violate the minimum security distance ($\kappa=2\text{ m.}$. in the experiments), potentially leading to collisions, which are notably present in Diffusion and CMF. To ensure a fair comparison, we have selected the most similar shapes generated by each method.
@@ 6 Conclusions
Acknowledgment-. This work was supported by a DGA scholarship and project T45_23R; MCIN/AEI/ ERDF/NextGenerationEU/PRTR projects PID2021-125514NB-I00, PID2021-125209OB-I00 and TED2021-129410B-I00, and by ONR research grants N62909-24-1-2081, and N00014-23-1-2354.

### v2

@@ Abstract
Gen-Swarms is an innovative method that leverages and combines the capabilities of deep generative models with reactive navigation algorithms to automate the creation of drone shows. Recent advancements in deep generative models, particularly diffusion models, have demonstrated remarkable effectiveness in generating high-quality 2D images. Building on this success, various works have extended diffusion models to 3D point cloud generation. In contrast, alternative generative models such as flow matching have been proposed, offering a simple and intuitive transition from noise to meaningful outputs. However, the application of flow matching models to 3D point cloud generation remains largely unexplored. Gen-Swarms adapts these models to automatically generate drone shows. Existing 3D point cloud generative models create point trajectories which are impractical for drone swarms. In contrast, our method not only generates accurate 3D shapes but also guides the swarm motion, producing smooth trajectories and accounting for potential collisions through a reactive navigation algorithm incorporated into the sampling process. For example, when given a text category like Airplane, Gen-Swarms can rapidly and continuously generate numerous variations of 3D airplane shapes. Our experiments demonstrate that this approach is particularly well-suited for drone shows, providing feasible trajectories, creating representative final shapes, and significantly enhancing the overall performance of drone show generation. Code has been released to the community  [^1] Homepage located at: https://cplou99.github.io/Gen-Swarms/.
@@ 3.2 DDPMs and CNFs
$\mathbf{x}_{t}=\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}+\sqrt{\beta_{t}}\epsilon,\;\;\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\;\;\;t\in\{1,\ldots,T\},$ (1)
@@ 3.3 Flow matching for generative modeling
$p_{t}(\mathbf{x}|\mathbf{x}_{0}):=\mathcal{N}\left(\mathbf{x}|t^{\prime}\mathbf{x}_{0},\sigma_{t}\mathbf{I}\right),\;\;\sigma_{t}=1-(1-\sigma_{\min})t^{\prime},\;\;t^{\prime}=(T-t)/T,$ (4)
@@ 4.1 Method
$p_{t}(\mathbf{X}|\mathbf{X}_{0}):=\mathcal{N}\left(\mathbf{X}|t^{\prime}\mathbf{X}_{0},\sigma_{t}\mathbf{I}\right),\;\;\sigma_{t}=1-(1-\sigma_{\min})t^{\prime},\;\;t^{\prime}=(T-t)/T.$ (8)
@@
$q(\mathbf{z})=q_{\mathbf{w}}\left(\mathbf{w}\right)\bigg{|}\frac{\partial F_{\alpha}}{\partial\mathbf{w}}\bigg{|}^{-1},\;\;\mathbf{w}=F^{-1}_{\alpha}(\mathbf{z}).$ (10)
@@
$\mathbf{X}_{t-\Delta t}=\mathbf{X}_{t}+\rho(\mathbf{V}^{t}_{\theta},\mathbf{X}_{t},\kappa),\;\;\;\;\forall t\in\{T,T-\Delta t,\ldots,\Delta t\}.$ (13)
@@ 5.2 Experimental Setup
Regarding architecture details, our Gated Contextual Network consists of $6$ layers, while the bijector $F_{\alpha}$ comprises 14 coupling layers. These neural networks, along with the PointNet encoder with embedding size $d=256$, are trained using the Adam optimizer with a linear scheduler for the learning rate. The experiments were conducted on an Intel Core™ i7-12700K processor with 20 cores and an NVIDIA GeForce RTX 3090 GPU. Furthermore, focusing on method hyperparameters, we set $T=1$ and vary the $\Delta t$ rate according to the experiment to ensure a fair comparison. Furthermore, we set a minimum security distance between points of $\kappa=2$ meters in our final scale, resulting in $\kappa=0.06$ meters on the training scale. This distance is based on the assumption that drones are 0.5 meters in size (0.25m of radius) with a security distance of 1.5 meters between drones. Lastly, we choose ORCA [30] as our reactive navigation method $\rho$.
@@
  • CFM+ORCA: This third baseline is designed to assess the performance of Gen-Swarms sampling. This approach takes the final point cloud generated by CFM and uses directly the ORCA reactive navigation method to generate all trajectories from the random initial point cloud towards the goal.
@@ 5.3 Quantitative results
For each method, we generate 607 samples, matching the test set size to obtain accurate quality metrics for 3D shapes. Additionally, to ensure a fair comparison, we perform sampling over 100 steps ($\Delta t=0.01s.$) for all methods. We compare the performance of Gen-Swarms against the baselines in the Airplane category using the metrics outlined in Sec. 5.2. The results, presented in a real scale of $200\times 200\times 200$ meters, are summarized in Table 1. As observed, Diffusion leads in quality but falls short in collision avoidance, smoothness and energy consumption. Conditional Matching Flow improves the smoothness and provides more direct, lower-energy trajectories, but still suffers from collisions. When leveraging ORCA to generate safe trajectories towards the final 3D point clouds generated by CFM, it fails to reach these positions within the same number of steps, and does not form the final desired shape, as graphically visualized in the fourth column of Fig. 5. In contrast, Gen-Swarms delivers balanced performance, excelling in collision avoidance (collisions are only present in the initial random point cloud, which is why the TRAJ metric is not zero), energy efficiency, and maintaining high quality while ensuring smooth trajectories. This is because Gen-Swarms uses CFM as a closed-loop controller with ORCA as an additional reactive navigation layer inside the loop.
@@ 5.4 Qualitative results
We illustrate in Fig. 5 an example of a final 3D shape produced by each method in the Airplane category. Drones highlighted in red would violate the minimum security distance ($\kappa=2\text{ m.}$. in the experiments), potentially leading to collisions, which are notably present in Diffusion and CFM. To ensure a fair comparison, we have selected the most similar shapes generated by each method.
@@ 6 Conclusions
Acknowledgments
This work was supported by a DGA scholarship and project T45_23R; MCIN/AEI/ ERDF/NextGenerationEU/PRTR projects PID2021-125514NB-I00, PID2021-125209OB-I00 and TED2021-129410B-I00, and by ONR research grants N62909-24-1-2081, and N00014-23-1-2354.